OpenStudy (anonymous):

Find the equation of a line that contains (3,4) and has a slope of 2. Write in point-slope form.

OpenStudy (anonymous):

Hint: Point-Slope form y - y1 = m(x - x1) m = slope Substitute the x and y-values from the given point into the equation.

OpenStudy (anonymous):

Does that help?